OpenStudy (anonymous):

Find the 7th term in the geometric sequence where t1 = 10 and r = -3.

sam (.sam.):

\[\large T_n=a_1r^{n-1} \\ \\ \large T_7=(10)(-3)^{7-1}=?\]

OpenStudy (anonymous):

(10)(-3^7-1

OpenStudy (anonymous):

-30

OpenStudy (anonymous):

?

sam (.sam.):

Erm\[T_7=(10)(-3)^{6} \\ \\ T_7=(10)(729)=?\]

OpenStudy (anonymous):

7290

sam (.sam.):

yes \(\large \color{blue}{\checkmark}\)

OpenStudy (anonymous):

and....

sam (.sam.):

?

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

got 1 more question though

sam (.sam.):

ok